OpenStudy (ria23):

y=x^2-2x-3 Would the vertex and intercept form be: y+4=(x-1)^2 y=(x+3)(x-1) ?

jimthompson5910 (jim_thompson5910):

y+4=(x-1)^2 is correct

jimthompson5910 (jim_thompson5910):

but y=(x+3)(x-1) is incorrect

jimthompson5910 (jim_thompson5910):

notice how the +3 and the -1 in y=(x+3)(x-1) add to +3 + (-1) = +2 but we want them to add to -2 (for the -2x)

OpenStudy (ria23):

So, (x-3)(x+1)?

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (ria23):

^.^ Thank yhu

jimthompson5910 (jim_thompson5910):

np